Dear Tutors can you please assist me with this problem thank you
A line with a slope of -10 passes through the points (8,-5) and (h,5). What is the value of h?
h=

Answer by Theo(13342)

slope is equal to (y2 - y1) / (x2 - x1)

(x1,y1) is a point on the line.

(x2,y2) is another point on the line.

assign (8,-5) to (x1,y1)

assign (h,5) to (x2,y2)

slope formula becomes (5 - (-5)) / (h - 8)

simplify to get slope = 10 / (h-8)

since slope = -10, the equation becomes -10 = 10 / (h-8)

multiply both sides of the equation by (h-8) to get -10 * (h - 8) = 10

simplify to get -10h + 80 = 10

subtract 80 from both sides of the equation to get -10h = -70

divide both sides of the equation by -10 to get h = 7

the two points are:

(x1,y1) = (8,-5)

(x2,y2) = (7,5)

(y2-y1) / (x2-x1) = 10/-1 = -10

solution looks good.

h = 7